OpenStudy (anonymous):

(tan(180-a)(tan(270-a)). Find product.

OpenStudy (bradely):

(tan(180-a)(tan(270-a)) Here,tan(180-a) =-tan a tan(270-a) =cota (tan(180-a)(tan(270-a))=-tana*cota =-1 Source: http://www.mathskey.com/question2answer/ask